Circular No. 8163

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(5381) SEKHMET M. C. Nolan and E. S. Howell, National Astronomy and Ionosphere

Center; A. S. Rivkin, Massachusetts Institute of Technology; and C. D. Neish, University of British Columbia, write: “Arecibo radar delay-Doppler images (2380 MHz, 12.6 cm) obtained on May 8–11 show that minor-planet (5381) is a binary system. Preliminary estimates of average diameters, based on range extents at 45-m resolution, are 1000 m and 300 m. The orbital period appears to be near 12 hr, with 24 hr much less probable. The minimum orbital radius is 1.3 km, with 1.5 km as our best estimate. Visible spectroscopy obtained on May 21 at the Kitt Peak 4-m telescope show an absorption band near 1 μm, suggesting a stony composition. Optical lightcurve observations of this object would be extremely valuable.”

SUPERNOVA 2003gl IN NGC 7782 Further to IAUC 8162, M. Schwartz and P. Holvorcem report the dis-

covery via LOTOSS of an apparent supernova on unfiltered Tenagra II images taken on July 5.4 (mag ∼ 18.3) and 6.4 UT (mag ∼ 18.2). SN 2003gl is located at α = 23^{h}53^{m}53^{s}.99, δ = +7^{o}57′23′′.7 (equinox 2000.0), which is 1′′.9 east and 51′′.2 south of the nucleus of NGC 7782. A Tenagra II image taken on 2002 Nov. 23.3 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2003gd IN M74 P. Garnavich, University of Notre Dame; and E. Bass, Cornell Uni-

versity, report that the position of SN 2003gd was included on R-band images taken with the Wisconsin-Indiana-Yale-NOAO telescope (WIYN + MiniMosaic Camera) on 2003 Jan. 29.1 UT, while observing SN 2002ap (total exposure time 900 s; 1′′.0 seeing). Using the accurate astrometry of McNaught (IAUC 8152) and comparing with the Hubble Space Telescope archival images (Smartt et al., IAUC 8152), the WIYN images place a limit of R > 24 on the progenitor brightness at four months before discovery. This observation, combined with spectral-age estimates by Kotak et al. and Phillips, constrains the likely date of explosion to between February and April.

2003 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green